The area of a rectangle is 234 square feet. If the perimeter is 62 feet find the length o the rectangle
Answer by checkley77(12844)

2L+2W=62
2L=62-2W
L=(62-2W)/2
L=31-W
LW=234
(31-W)W=234
31W-W^2-234=0
-W^2+31W-234=0
W^2-31W+234=0
(W-18)(W-13)=0
W-18=0
W=18 ANS L=13 ANS.
W-13=0
W=13 ANS L=18 ANS.
PROOF:
2*18+2*13=62
36+39=62
62=62
13*18=234
234=234